# Rectifying Conformity Scores for Better Conditional Coverage

### v1

@@ 5 Related works
A first class of methods involves group-conditional guarantees (Ding et al., 2024; Jung et al., 2023), which provide coverage guarantees for a predefined set of groups. Another class partitions the covariate space into multiple regions and applies classical conformal prediction within each region (LeRoy & Zhao, 2021; Alaa et al., 2023; Kiyani et al., 2024). The significant limitation of these methods lies in the need to specify the groups or regions in advance.
@@
Methods utilizing conditional density estimation have been proposed to produce conformal prediction intervals that adapt to skewed data (Sesia & Romano, 2021), to minimize the average volume (Sadinle et al., 2019, denoted DCP in our paper) or to define more flexible highest-density regions (Izbicki et al., 2022; Plassier et al., 2024a). Probabilistic conformal prediction (PCP; Wang et al., 2023) bypasses density estimation by constructing prediction sets as unions of balls centered on samples from a generative model. All these methods are either tailored to handle the scalar response ($d=1$) or require an accurate conditional distribution estimate which might be hard to obtain in practical scenarios.
@@
Guan (2023) introduces a localized conformal prediction framework that adapts to data heterogeneity by weighting calibration points based on their similarity to the test sample. To do so, kernel-based localizers assign greater importance to nearby points, tailoring prediction intervals to local data patterns. Amoukou & Brunel (2023) extend Guan’s approach by replacing kernels with quantile regression forest estimators for improved performance. Although effective, these methods face challenges in high-dimensional or mixed-variable settings. Here again, these methods have mostly been used in the case where $d=1$.
@@
Several methods aim to transform conformity scores to improve approximate conditional coverage. For example, Han et al. (2022) presents an approach that uses kernel density estimation to approximate the conditional distribution. Similarly, Deutschmann et al. (2023) rescales the conformity scores based on an estimate of the local score distribution using the jackknife+ technique. However, these methods generally rely on estimating the conditional distribution of conformity scores, which is challenging in practice.
@@
Finally, various recalibration methods have been proposed to improve marginal coverage (Dheur & Taieb, 2023) or conditional coverage (Dey et al., 2022). While these methods can also be interpreted within the conformal prediction framework (Dheur & Taieb, 2024; Marx et al., 2022), they often require modifications to the training procedure, making them less broadly applicable than purely conformal methods.
@@ 6 Theoretical guarantees
$\mathbb{P}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})\,|\,X_{n+1}}\right)\geq 1-\alpha\\ +\epsilon_{\tau}(X_{n+1})-\alpha\mathrm{L}\times[{F_{\tilde{V}}(\varphi)}]^{n+1}.$ (17)
The proof is postponed to Section B.5. According to Theorem 4, the conditional validity of the prediction set $\mathcal{C}_{\alpha}(x)$ directly depends on the accuracy of the quantile estimate $\tau(x)$. If $\tau(x)$ closely approximates the conditional quantile $Q_{1-\alpha}\bigl{(}\textup{P}_{V_{\varphi}(x,Y)\mid X=x}\bigr{)}$, then (17) ensures that conditional coverage is approximately achieved.
@@ 7.1 Toy example
We set $\alpha=0.1$ and consider the conformity score $V(x,y)=|y-\mu(x)|$. In this case, the $(1-\alpha)$-quantile of $V(x,Y)\mid X=x$ is known and we denote it by $q_{1-\alpha}(x)$. Given $\omega\in[0,1]$, we set $\widehat{\tau}(x)\sim(1-\omega)q_{1-\alpha}(x)+\omega\epsilon(x)$, where we consider $\epsilon(x)\sim\mathcal{N}(0,x^{4})$. We perform $1000$ experiments and report the $10\%$ lower value of $x\in[0,1]\mapsto\mathbb{P}(Y_{n+1}\in\mathcal{C}_{\alpha}(x)\mid X=x)$; the results can be found in Table 1. If $\omega=0$, $\widehat{\tau}(x)$ corresponds to the true $(1-\alpha)$ quantile. In this case, our method is conditionally valid, as Theorem 4 shows. However, while all settings of $\omega$ yield marginally valid prediction sets, the conditional coverage decreases as the quantile estimate $\widehat{\tau}(x)$ deteriorates.
$\omega$ 00 $1/3$ $2/3$ $1$
@@ 7.2 Real-world experiment
where $\mathcal{N}\bigl{(}y\mid\mu_{k}(x),\Sigma_{k}(x)\bigr{)}$ is a Gaussian density with mean $\mu_{k}(x)$ and covariance $\Sigma_{k}(x)$.
Methods.  We compare RCP with four split-conformal prediction methods from the literature: ResCP (Diquigiovanni et al., 2024), PCP (Wang et al., 2023), DCP (Sadinle et al., 2019), and SLCP (Han et al., 2022). ResCP uses residuals as conformity scores. To handle multi-dimensional outputs, we follow (Diquigiovanni et al., 2024) and define the conformity score as the $l^{\infty}$ norm of the residuals across dimensions, i.e., $V(x,y)=\max_{i\in[d]}|\hat{\mu}_{i}(x)-y_{i}|$. PCP constructs the prediction set as a union of balls, while DCP defines the prediction set by thresholding the density. ResCP is compatible with the mean predictor, whereas PCP and DCP are compatible with the mixture predictor. Finally, SLCP, like RCP, is compatible with any conformity score and base predictor. For RCP, we compuate an estimate $\widehat{\tau}(x)$ (see Algorithm 1) using quantile regression with a fully connected neural network composed of 3 layers with 100 units.
@@
Evaluation metrics.  To evaluate conditional coverage, we use worst-slab coverage (Cauchois et al., 2020; Romano et al., 2020) and the conditional coverage error, computed over a partition of $\mathcal{X}$, following Dheur et al. (2024).
Figure 3: Worst-slab coverage for three conformal methods and their RCP counterparts, on datasets sorted by total size.
@@
Results.Figure 3 presents the worst-slab coverage for different conformity scores, both with and without RCP. Similarly, Figure 4 compares worst-slab coverage between SLCP and RCP. Additional results, including conditional coverage error and marginal coverage, are provided in Section A.1.
In Figure 3, we observe that ResCP, PCP, and DCP fail to reach the nominal level of conditional coverage for most datasets. In contrast, all variants of RCP significantly improve coverage across all datasets. Similarly, Figure 4 shows that RCP often achieves better conditional coverage than SLCP, particularly on larger datasets. Figure 5 in Section A.1 confirms these findings with the conditional coverage error. Finally, as expected, all methods achieve marginal coverage.
@@
Section A.5 presents an additional study comparing RCP with CP and CQR methods, that we adapted to multidimensional target setting. For these experiments, the model predicts parameters of a multivariate normal distribution and we use the score based on the corresponding Mahalanobis distance. We demonstrate that RCP improves conditional coverage over classic CP and also benefits from the custom score to outperform CQR.
@@ Appendix A Additional experiments
A.1 Additional results
Figure 5 extends the results in Figure 3 by displaying additionally the marginal coverage and conditional coverage error. As expected, all methods obtain a correct marginal coverage. Furthermore, the methods with the best worst slab coverage (closest to $1-\alpha$) also obtain a small conditional coverage error, supporting our conclusions in Section 7.
@@ A.3 Choice of adjustment function
Figure 7 compares RCP with difference (--) and linear (*) adjustments when combined with the DCP method. Since RCP with any adjustment function adheres to the SCP framework, marginal coverage is guaranteed, as shown in Panel 1.
The conformity score for DCP is defined as $V(x,y)=-\log\hat{p}(y\mid x)$, which can take negative values, implying that $\mathbb{T}=\mathbb{R}$. However, the linear adjustment requires $\mathbb{T}\subseteq\mathbb{R}_{+}^{*}$, violating 1 and resulting in a failure to approximate conditional coverage accurately. This issue is evident in Panel 2, particularly for datasets rf_1, rf_2, meps_21, meps_19, and meps_20. In contrast, the difference adjustment does not impose such a restriction.
@@ A.4 Additional adjustment functions
A.5 Comparison with simple baselines
Figure 10 demonstrates improvements of RCP over simpler baseline methods, serving as a small ablation study.
The first baseline, denoted CQR, is based on quantile regression estimates for each dimension of the output variable. First, (univariate) conformalized quantile regression scores (Romano et al., 2019) are computed for each dimension. Then, they are aggregated by taking the maximum score over each dimension, similarly to ResCP in the main text. The resulting prediction set in this case is a hyperrectangle. Its size is adaptive to the input, but the conformal correction is isotropic and constant for all input points.
For the remaining methods, the model outputs parameters of a multivariate normal distribution. The score in this case is the Mahalanobis distance. The first part of the name before the dash states the type of the base model. “Const” is a constant prediction of the empirical mean and covariance matrix on the train set. “MLP” corresponds to a small neural network that predicts these parameters based on the input point $x$. The prediction sets for this group of methods are hyperellipsoids (not necessarily aligned with the axes).

### v2

@@ 5 Related Work
A first class of methods involves group-conditional guarantees (Jung et al., 2023; Ding et al., 2024), which provide coverage guarantees for a predefined set of groups. Another class partitions the covariate space into multiple regions and applies classical conformal prediction within each region (LeRoy & Zhao, 2021; Alaa et al., 2023; Kiyani et al., 2024). The significant limitation of these methods lies in the need to specify the groups or regions in advance.
@@
Methods utilizing conditional density estimation have been proposed to produce conformal prediction intervals that adapt to skewed data (Sesia & Romano, 2021), to minimize the average volume (Sadinle et al., 2019, denoted DCP in our paper) or to define more flexible highest-density regions (Izbicki et al., 2022; Plassier et al., 2025). Probabilistic conformal prediction (PCP; Wang et al., 2023) bypasses density estimation by constructing prediction sets as unions of balls centered on samples from a generative model. All these methods are either tailored to handle the scalar response ($d=1$) or require an accurate conditional distribution estimate which might be hard to obtain in practical scenarios.
@@
Guan (2023) introduces a localized conformal prediction framework that adapts to data heterogeneity by weighting calibration points based on their similarity to the test sample. To do so, kernel-based localizers assign greater importance to nearby points, tailoring prediction intervals to local data patterns. Amoukou & Brunel (2023) extend Guan’s approach by replacing kernels with quantile regression forest estimators for improved performance. Although effective, these methods face challenges in high-dimensional or mixed-variable settings.
@@
Several methods aim to transform conformity scores to improve approximate conditional coverage. For example, Johansson et al. (2021), following earlier works by Papadopoulos & Haralambous (2011); Johansson et al. (2014); Lei et al. (2018), investigate normalized conformity scores (NCF), which enhance standard conformal prediction by adjusting prediction set according to instance difficulty. NCF can be represented within our framework through a specific choice of the function $f_{\tau}(v)=v/(\tau+\beta)$, where $\beta$-values will put a greater emphasis on the difficulty estimation. Notably, the estimation approach employed in these papers uses least-squares regression on residuals, in contrast to the quantile regression approach adopted in RCP, which is essential to satisfy (4). Han et al. (2022) presents an approach that uses kernel density estimation to approximate the conditional distribution. Similarly, Deutschmann et al. (2023) rescales the conformity scores based on an estimate of the local score distribution using the jackknife+ technique. However, these methods generally rely on estimating the conditional distribution of conformity scores, which is challenging in practice. Dewolf et al. (2025) studies conditional validity of normalized conformal predictors in oracle setting, i.e., when the optimal normalization is known.
Recent work by Colombo (2024) suggests to transform the conformity score employing a normalizing flow: $\tilde{V}(x,y)=b(V(x,y),x)$. The normalizing flow is trained to map the joint distribution $\textup{P}_{V,X}$ of the conformity score and attributes into a product distribution, $\textup{P}_{\tilde{\mathbf{V}}}\otimes\textup{P}_{X}$, where $\textup{P}_{\tilde{\mathbf{V}}}$ is an arbitrary univariate distribution. Notably, this condition is stricter than the conditional coverage criterion (4), as it enforces $\textup{P}_{\tilde{\mathbf{V}}|X=x}=\textup{P}_{\tilde{\mathbf{V}}}$ for almost every $x$ under $\textup{P}_{X}$. Consequently, learning such a transformation typically necessitates a larger sample size; see Section 7.
One method (Xie et al., 2024) proposes to use a cross-validated boosting procedure to learn a new score function to be used in split-conformal prediction. The authors consider a specific family of possible score functions and corresponding loss functions tailored either to deviation from conditional coverage or interval length. This method has several limitations compared to our approach: limited set of score functions, tailored to one-dimensional targets, requires access to the train set, and high computation cost.
RCP method shares some similarities with that of Gibbs et al. (2025), which also performs a quantile regression of the conformity score with respect to the attribute $X$. There are two essential differences: firstly, Gibbs et al. (2025) work directly with the conformity score $V$, whereas we regress on a transformed score $V_{\varphi}$. Secondly, the manner in which the quantile regression result is used differs significantly. RCP uses the quantile estimator to define the rectified score $\tilde{V}$, to which the standard CP procedure, while Gibbs et al. (2025) propose a considerably more complex procedure; see Section 7.
@@
Finally, various recalibration methods have been proposed to improve marginal coverage (Dheur & Taieb, 2023) or conditional coverage (Dey et al., 2022). While these methods can also be interpreted within the conformal prediction framework (Marx et al., 2022; Dheur & Taieb, 2024), they often require modifications to the training procedure, making them less broadly applicable than purely conformal methods.
@@ 6 Theoretical Guarantees
$\mathbb{P}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})\,|\,X_{n+1}=x}\right)\geq 1-\alpha\\ +\epsilon_{\hat{\tau}}(x)-\alpha\mathrm{L}\times[{F_{\tilde{V}}(\varphi)}]^{n+1}.$ (16)
The proof is postponed to Section B.5. According to Theorem 2, the conditional validity of the prediction set $\mathcal{C}_{\alpha}(x)$ directly depends on the accuracy of the quantile estimate $\widehat{\tau}(x)$. If $\widehat{\tau}(x)$ closely approximates the conditional quantile $Q_{1-\alpha}\bigl{(}\textup{P}_{V_{\varphi}(x,Y)\mid X=x}\bigr{)}$, then (16) ensures that conditional coverage is approximately achieved.
@@ 7.1 Toy example
We set $\alpha=0.1$ and consider the conformity score $V(x,y)=|y-\mu(x)|$. In this case, the $(1-\alpha)$-quantile of $V(x,Y)\mid X=x$ is known and we denote it by $Q_{1-\alpha}(x)$. Given $\omega\in[0,1]$, we set $\widehat{\tau}(x)\sim(1-\omega)Q_{1-\alpha}(x)+\omega\epsilon(x)$, where we consider $\epsilon(x)\sim\mathcal{N}(0,x^{4})$. We perform $1000$ experiments and report the $10\%$ lower value of $x\in[0,1]\mapsto\mathbb{P}(Y_{n+1}\in\mathcal{C}_{\alpha}(x)\mid X=x)$; the results can be found in Table 1. If $\omega=0$, $\widehat{\tau}(x)$ corresponds to the true $(1-\alpha)$-quantile. In this case, our method is conditionally valid, as Theorem 2 shows. However, while all settings of $\omega$ yield marginally valid prediction sets, the conditional coverage decreases as the quantile estimate $\widehat{\tau}(x)$ deteriorates.
$\omega$ 0 $1/3$ $2/3$ 11
@@ 7.2 Real-world experiment
where $\mathcal{N}\bigl{(}y\mid\mu_{k}(x),\Sigma_{k}(x)\bigr{)}$ is a Gaussian density with mean $\mu_{k}(x)$ and covariance matrix $\Sigma_{k}(x)$.
Methods.  We compare RCP with four split-conformal prediction methods from the literature: ResCP (Diquigiovanni et al., 2024), PCP (Wang et al., 2023), DCP (Sadinle et al., 2019), and SLCP (Han et al., 2022). ResCP uses residuals as conformity scores. To handle multi-dimensional outputs, we follow (Diquigiovanni et al., 2024) and define the conformity score as the $l^{\infty}$ norm of the residuals across dimensions, i.e., $V(x,y)=\max_{i\in[d]}|\hat{\mu}_{i}(x)-y_{i}|$. PCP constructs the prediction set as a union of balls, while DCP defines the prediction set by thresholding the density. ResCP is compatible with the mean predictor, whereas PCP and DCP are compatible with the mixture predictor. Finally, SLCP, like RCP, is compatible with any conformity score and base predictor. For RCP, we compute an estimate $\widehat{\tau}(x)$ (see Algorithm 1) using quantile regression with a fully connected neural network composed of 3 layers with 100 units.
@@
Figure 3: Worst-slab coverage and volume for three conformal methods and their RCP counterparts, on datasets sorted by total size.
Evaluation metrics.  To evaluate conditional coverage, we use worst-slab coverage (WSC, Cauchois et al., 2020; Romano et al., 2020) with $\delta=0.2$ and the conditional coverage error, computed over a partition of $\mathcal{X}$, following Dheur et al. (2025). To evaluate sharpness, we also report the median of the logarithm of the prediction set volume, scaled by the dimension $d$.
Main results.Figure 3 presents the worst-slab coverage and volume for different conformity scores, both with and without RCP. Similarly, Figure 4 compares worst-slab coverage between SLCP and RCP. Additional results, including conditional coverage error and marginal coverage, are provided in Section A.1.
@@
In the left panel of Figure 3, we observe that ResCP, PCP, and DCP fail to reach the nominal level of conditional coverage for most datasets. In contrast, all variants of RCP significantly improve coverage across all datasets. Similarly, Figure 4 shows that RCP often achieves better conditional coverage than SLCP, particularly on larger datasets. Figure 5 in Section A.1 confirms these findings with the conditional coverage error. Finally, as expected, all methods achieve marginal coverage.
In the right panel of Figure 3, we observe that RCP improves the median prediction set volume compared to non-RCP variants in addition of improving conditional coverage.
Finally, Section A.6 compares average and median volumes of prediction sets produced by direct conformal methods and their RCP counterparts. Direct methods obtain a smaller average volume while RCP obtains a smaller median volume.
Additional experiments.  We complement main results with multiple experiments aiming at studying variations of the proposed method and comparing it with some additional competitors.
@@
Section A.5 directly compares the proposed method with CQR, showing that CQR already obtains a competitive conditional coverage but is outperformed by RCP-DCP in average volume. Section A.7 presents an additional study comparing RCP with CP and CQR methods, that we adapted to multidimensional target setting. For these experiments, the model predicts parameters of a multivariate normal distribution and we use the score based on the corresponding Mahalanobis distance. We demonstrate that RCP improves conditional coverage over classic CP and also benefits from the custom score to outperform CQR.
Section A.8 considers an approach to improve data efficiency. Instead of dividing the calibration dataset $\mathcal{D}$ into two parts to estimate $\widehat{\tau}$, we compute out-of-sample conformity scores on the training dataset $\mathcal{D}_{\text{train}}$ using $K$-fold cross-validation. This results in improved conditional coverage at the cost of training $K$ additional models.
Section A.9 provides an additional comparison with Conditional Prediction with Conditional Guarantees (CPCG; Gibbs et al. (2025)). CPCG obtains a competitive conditional coverage but is 200-100000 times slower than RCP overall, limiting its applicability.
@@ 8 Conclusion
Acknowledgments
V.P. carried out this study during his PhD under the supervision of E.M., and subsequently as a research engineer at the Centre Lagrange in Paris. Part of E.M.’s work was also conducted under the auspices of the Centre Lagrange. E.M. was all partially funded by the European Union (ERC-2022-SyG, 101071601). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Impact Statement
This work contributes to the broader effort to improve the interpretability and statistical reliability of machine learning algorithms. Prediction intervals with exact marginal and approximate conditional coverage offer a useful tool for conveying uncertainty regarding the accuracy of ML models, which is essential for increasing their fairness, and fostering wider acceptance.
@@ Appendix A Additional Experiments
A.1 Additional results on marginal coverage and conditional coverage error
Figure 5 extends the results of Figure 3 by displaying additionally the marginal coverage and conditional coverage error. As expected, all methods obtain a correct marginal coverage. Furthermore, the methods with the best worst slab coverage (closest to $1-\alpha$) also obtain a small conditional coverage error, supporting our conclusions in Section 7.
@@ A.3 Choice of adjustment function
Figure 7 compares RCP with difference (-) and linear (*) adjustments when combined with the DCP method. Since RCP with any adjustment function adheres to the SCP framework, marginal coverage is guaranteed, as shown in Panel 1.
The conformity score for DCP is defined as $V(x,y)=-\log\hat{p}(y\mid x)$, which can take negative values, implying that $\mathbb{T}=\mathbb{R}$. However, the linear adjustment requires $\mathbb{T}\subseteq\mathbb{R}_{+}^{*}$, violating 1 and resulting in a failure to approximate conditional coverage accurately. This issue is evident in Panel 2. In contrast, the difference adjustment does not impose such a restriction.
@@ A.4 Additional adjustment functions
A.5 Direct comparison with CQR
Here we present a direct comparison of RCP with Conformalized Quantile Regression (CQR; Romano et al. (2019)). We use the same underlying neural network architectures for the models as in our main experiment. Similarly to ResCP, to handle multi-dimensional outputs, we follow (Diquigiovanni et al., 2024) and define the conformity score of CQR as the $l^{\infty}$ norm of the CQR conformity scores across dimensions. Specifically, we compare CQR to DCP and its RCP-DCP counterpart, which achieves the best median volume overall.
Figure 10 shows that CQR matches the conditional coverage of RCP-DCP. However, it produces larger median prediction sets due to less flexible shapes.
Figure 10: Worst slab coverage and (logarithm) median prediction set volume (scaled by $d$).
A.6 Comparison of prediction set volumes
Table 2 shows the average volume obtained by the methods compared in Section 7. Non-RCP methods obtain a smaller average volume across all datasets. The larger average volume of RCP is explained by the larger regions produced for instances with larger uncertainty.
Table 2: Mean prediction set volume per dataset.
dataset PCP RCP-PCP DCP RCP-DCP ResCP RCP-ResCP
households 88.3 1.33e+02 47.4 1.02e+02 1.81e+02 4.51e+02
scm20d 4.26e+05 7.92e+06 1.11e+06 3.95e+07 5.22e+05 7.15e+12
rf1 0.0274 0.190 0.000562 4.35e+04 0.0276 7.94e+08
scm1d 1.92e+04 1.30e+08 2.30e+04 1.67e+08 6.27e+04 2.04e+15
meps_21 1.65 10.0 0.746 6.07 5.35 8.32e+12
meps_19 90.0 3.27e+04 3.64 3.27e+04 5.56 3.56e+22
meps_20 1.68 5.50 0.761 6.20 5.38 6.27e+13
house 0.676 0.936 0.519 0.751 2.92 3.88
bio 0.579 1.12 0.414 0.645 1.05 1.16
blog_data 0.459 8.45e+02 0.143 6.37e+02 1.26 6.79e+21
calcofi 3.47 4.12 2.45 3.06 4.53 4.47
taxi 9.21 9.63 5.69 6.40 12.4 12.8
In contrast, Table 3 shows that RCP obtains smaller regions across most datasets when comparing the median volume, avoiding outliers.
Table 3: Median prediction set volume per dataset.
dataset PCP RCP-PCP DCP RCP-DCP ResCP RCP-ResCP
households 67.4 56.5 39.2 32.1 1.81e+02 1.67e+02
scm20d 3.11e+04 1.42e+04 7.03e+05 7.74e+04 5.22e+05 2.00e+05
rf1 0.0110 0.00525 0.000583 33.1 0.0276 0.0231
scm1d 1.16e+02 2.05 1.01e+04 25.7 6.27e+04 1.70e+03
meps_21 1.04 0.704 0.433 0.227 5.35 2.42
meps_19 3.85 0.754 2.10 0.303 5.56 2.54
meps_20 1.05 0.616 0.416 0.254 5.38 2.51
house 0.596 0.519 0.471 0.386 2.92 2.67
bio 0.507 0.435 0.374 0.344 1.05 0.829
blog_data 0.229 0.209 0.0869 0.0597 1.26 1.16
calcofi 3.83 3.98 2.85 2.77 4.53 4.75
taxi 8.67 8.25 5.22 5.27 12.4 10.1
A.7 Case of ellipsoidal prediction sets
In this section, we will investigate how all parts of our proposed RCP method contribute to the performance. Additionally, we demonstrate the wider applicability of our approach: in this section, we use a different score and ellipsoid prediction sets. To achieve this, we modify our base models to predict parameters of the multivariate normal distribution. As a baseline method, we have selected CQR because of its popularity and ease of use.
Figure 11 demonstrates improvements of RCP over simpler methods, with CQR serving as a strong contemporary alternative. Each of these simpler methods employs a consecutively more complex model and/or calibration procedure. The first part of the name corresponds to the base prediction model, and the second part after the dash denotes the calibration procedure:
  • Const: the base model for these methods is a constant prediction of multivariate normal distribution parameters estimated on the train set.
  • MLP: the base model is a multidimensional perceptron that predicts the parameters of multivariate normal distribution.
  • CP: classic conformal prediction using full calibration set. We estimate a fixed prediction ellipsoid size using Mahalanobis distance-based nonconformity score.
  • RCP: our usual RCP procedure where we split the calibration set and fit a quantile regression model to predict the $(1-\alpha)$ quantile of the Mahalanobis score. Similarly to the other experiments, quantile regression is fit using MLP underlying model.
The alternative method CQR is based on quantile regression estimates for each dimension of the output variable. First, (univariate) conformalized quantile regression scores (Romano et al., 2019) are computed for each dimension. Then, they are aggregated by taking the maximum score over each dimension, similarly to ResCP in the main text. The resulting prediction set in this case is a hyperrectangle. Its size is adaptive to the input, but the conformal correction is isotropic and constant for all input points.